An unknown charge is repelled by a force of 40 newtons when it is at a distance of 5.0 cm from a charge of 40 microcoulombs. Find the magnitude of the unknown charge.

1
Expert's answer
2022-03-20T18:57:55-0400

Electric force

"F=\frac{kq_1q_2}{r^2}"

"40=\frac{9\times10^9\times40\times10^{-6}\times q_2}{0.05^2}"

"q_2=2.77\times10^{-7}C"